# Specific heat of Gd3+ and Eu2+-based magnetic compounds

### v4

@@ VII Fits of experimental data
The quality of the fits is excellent. For comparison, a mean-field approach would fail to reproduce the specific heat curves, particularly the shape of the peak and the high-temperature tail (see MF specific heat in Figures 2 and 3).
@@
Table 4: Parameters used in the fits: $T_{N}$ is the Néel temperature, $z_{0}$ is the number of nearest neighbors in the reported structure (compatible with 3D connectivity), $z$ (fit) is the effective number of neighbors, $J$ is the interaction strength, and $K$ is the anisotropy. From the fit we can only conclude that $|J_{1}|<|J_{2}|$, see main text. This value correspond to $|J_{1}|=|J_{2}|$ and $|J_{3}|=0.05K$, see main text.
GdNi3Ga9 [34, 35] exhibits a layered honeycomb structure with $z_{0}=4$ (three intralayer and one interlayer nearest neighbors), consistent with $z=4$ found in the fit. The specific heat fit (Figure 8a) reproduces the experimental data with excellent accuracy. The low $z$ value indicates significant spatial and quantum fluctuations accounting for the observed broadened specific heat peak. As the LHc(4,AF) lattice (corresponding to an antiferromagnetic interaction) has $T_{N}/J_{1}\sim 13$ while experimentally $T_{N}=20K$ we can deduce $J\sim 20$$13\sim 1.5K$. This result of a dominant first-neighbor antiferromagnetic interaction is further supported by the observation of a small (1.8°) deviation from perfect AF alignment between magnetic moments in adjacent layers [35], which may arise from weak Dzyaloshinskii-Moriya interactions (DMI).
@@
GdCu2Ge2 has a tetragonal structure [36]. The Gd sublattice adopts a body-centered tetragonal structure. As this lattice is quite elongated ($c/a\gg 1$) if only the first four nearest-neighbor interactions were considered, the system would become two-dimensional with a broadened transition near $T_{N}$. The four next-nearest neighbors still lay in the same plane as the NN. To recover a tridimensional connectivity compatible with the experiments at least the eight following neighbors must be considered. This leaves as a minimum option to consider the first and third nearest neighbors ($z_{0}=12$) interactions. As explained before, this should be similar to consider other lattice with $z=12$ as the FCC lattice. In panel b) we show that indeed the specific heat can be described by any of these lattices. If, to simplify, we consider $J_{1}$ ($z_{1}=4$) equal to $J_{2}$ ($z_{2}=4$) we get the equation
@@
This assumption leads to three possibilities:
  • $J_{3}=0$ ($z=8$) which we can disregard has it implies a two dimensional behavior
@@
  • $|J_{3}/J_{1}|=3.7$ ($z=12$) which we can disregard has it seems unlikely to have a larger strength associated to a more distant bond.
  • $|J_{3}/J_{1}|=0.27$ ($z=12$) appears to be the only reasonable solution.
@@
Without the actual simulation for this particular combination of couplings, we can only estimate the coupling strengths To do so, we use the mean-field result for the critical temperature (see section V.1).
$\frac{k_{B}T_{C}}{\frac{S(S+1)}{3}}=(z_{1}+z_{2})|J_{1}|+z_{3}|J_{3}|=(8+8\times 0.27)|J_{1}|$
@@
from which we obtain $|J_{1}|=|J_{2}|\sim 0.20$ K and $|J_{3}|\sim 0.05$ K. From these results we can not determine the signs of the coupling, but from the antiferromagnetic nature of the compound, one infers that at least one of these couplings should be antiferromagnetic.
@@
GdPdBi adopts a half-Heusler structure with the Gd sublattice forming a face-centered cubic (FCC) lattice. Fitting with the layered honeycomb lattice (LHc(10), $z=10$) requires an unexpectedly large $J_{2}/J_{1}$ ratio of approximately $6.1$. To validate our approach, we compared our results with inelastic neutron scattering (INS) measurements on the related compound GdPtBi ($T_{N}=8.5$ K)[37, 38]. The LSWT analysis of GdPtBi[38] reported $J_{1}=0.14$ K, $J_{2}=0.27$ K, $J_{3}=0.049$ K, and $K=0.0098$ K, with $|J_{2}|/|J_{1}|>1$ (specifically $|J_{2}|/|J_{1}|\approx 1.93$), demonstrating that second-neighbor interactions dominate in this compound and that the anisotropy is small ($|K|/|J_{1}|\lesssim 0.1$). In GdPtBi, the magnetic structure consists of ferromagnetic (111) planes antiferromagnetically stacked along [111][37, 39], exhibiting low frustration. This structural and magnetic context validated the plausibility of the $|J_{2}|/|J_{1}|\gtrsim 1$ ratio for GdPdBi. But as shown in section VI the specific heat for $|J_{1}|<|J_{2}|$ is almost independent of $J_{2}$ and does not allow us to determine a ratio of interactions.
@@
The crystal structure of GdNiSi3 is an orthorhombic system. Gd atoms forms a bilayered arrangement. Only interaction at the 7th nearest neighbor connects bilayers to provide a tridimensional connectivity ($z_{0}=17$). A good description of the experimental data is obtain considering up to 10th nearest-neighbor ($z=25$) while the best fit is seen considering up to 13th NN ($z=33$). This suggests that long range interactions, well beyond first neighbors, are relevant in this compound. Given the complexity of these interactions we only consider the uniform $J_{\delta}$ case. The GdNiSi3(33) lattice has $T_{C}/{|J|}\sim 145$ while experimentally $T_{N}=22$K so we can deduce $|J|\sim 22$K/145=0.15/145=0.15K. The large amount of couplings needed (up to 13 nearest-neighbors) certainly come from a distribution of interactions. This distribution could be determined with first principle calculations that considers at least this coordination.
@@
This completes the Gd compounds that we have considered. In a previous work [33], the specific heat of GdCoIn5 has been fitted with essentially the same procedure. The compound has a tetragonal structure. The authors have extrapolated to the thermodynamic limit the specific heat that corresponds to a simple cubic lattice $z=6$ nearest-neighbor with antiferromagnetic interactions, providing an excellent fit of the data.
@@
The Eu sublattice of Eu2Pd3Sn can be though as a stacking of corrugated hexagonal layers in the ab plane. The smaller coordination number to have a tridimensional connectivity is $z_{0}=4$. This involves up to the third coordination shell of nearest neighbors (first and second nearest neighbors lie in the same layer). To properly compare the specific heat of Eu2Pd3Sn [41] with a Hamiltonian model we have to consider up to 6th nearest-neighbor in this lattice leading to a coordination $z=14$. The point group at the Eu sites is $C_{1h}$, which allows for the inclusion of the anisotropy term $H_{A}$, as discussed in section II.2. So an anisotropy term with $K/|J_{1}|=-0.4$ was included. With this anisotropy the Eu2Pd3Sn(14) lattice has $T_{C}/{|J_{0}|}\sim 66$. Experimentally $T_{N}=47$K so we can deduce, assuming uniform couplings, $|J|\sim 47$K/66=0.71/66=0.71K.
Finally, the Eu sublattice of EuPdSn2 can be though as a stacking of rectangular layers in the ac plane. The smaller coordination number to have a tridimensional connectivity is $z_{0}=8$. This involves up to the third coordination shell of nearest-neighbor interactions. To properly compare the specific heat of EuPdSn2 [29] with a Hamiltonian model we have to consider up to 7th nearest-neighbor in this lattice leading to a coordination $z=22$. The point group at the Eu sites is $C_{2v}$, which also allows for the inclusion of the term $H_{A}$. So an anisotropy term with $K/|J_{0}|=-0.9$ was included. With this anisotropy the EuPdSn2 lattice has $T_{C}/{|J_{0}|}\sim 112$. Experimentally $T_{N}=12$K from which we deduce $|J|\sim 12$K/112=0.11/112=0.11K. The discrepancies near $T_{N}$ likely originate from factors not explicitly included in our model, such as more complex magnetic ordering (e.g., a collinear amplitude-modulated structure) resulting from modulated interactions and crystal-field effects, as discussed in Ref. 14.
@@
Figure 8: Experimental specific heat $C/T$ versus $T/T_{N}$ for six 4f7 compounds (dots) and our theoretical fits (lines). $K=0$ for all Gd-based compounds (GdNi3Ga9, GdCu2Ge2, GdPbBi, GdNiSi3). Inset for GdCu2Ge2 show $z$ as a function of $J_{3}/J_{1}$ (black line) and the fitting $z=12$ value (horizontal red line). Experimental data sources: Refs. 34, 35, 42, 36, 27, 29, 41.
@@ VIII Summary
The effective number of neighbors $z$ quantifies the strength of spatial and quantum fluctuations, with larger $z$ values corresponding to reduced fluctuations and specific heat curves closer to mean-field predictions. Crucially, we find that the specific heat is remarkably insensitive to the detailed lattice structure for a given $z$ as long as the couplings do not strongly increase with distance. This enables a universal description of diverse magnetic systems.
@@ Acknowledgements.
IX Appendix

### v5

@@ VII Fits of experimental data
The quality of the fits is excellent. For comparison, a mean-field approach would fail to reproduce the specific heat curves, particularly the shape of the peak and the high-temperature tail (see mean-field specific heat in Figures 2 and 3). For a quantitative measure of the fits, see the discussion of Appendix E. The $L^{2}$ distance between the experimental curve and the corresponding fit is in all cases $\ell_{2}\leq 0.1$. Taking into account the experimental substraction of the phonon contribution, this is a reasonable distance.
@@
Table 4: Parameters used in the fits: $T_{N}$ is the Néel temperature, $z_{0}$ is the number of nearest neighbors in the reported structure (compatible with 3D connectivity), $z$ (fit) is the effective number of neighbors, $J$ is the interaction strength, and $K$ is the anisotropy. From the fit we can only conclude that $|J_{1}|<|J_{2}|$, see main text. This value correspond to $|J_{1}|=|J_{2}|=0.20K$ and $|J_{3}|=0.09K$, see main text.
GdNi3Ga9 [36, 37] exhibits a layered honeycomb structure with $z_{0}=4$ (three intralayer and one interlayer nearest neighbors), consistent with $z=4$ found in the fit. The specific heat fit (Figure 8a) reproduces the experimental data with excellent accuracy. The low $z$ value indicates significant thermal and quantum fluctuations accounting for the observed broadened specific heat peak. As the LHc(4,AF) lattice (corresponding to an antiferromagnetic interaction) has $T_{N}/J_{1}\sim 13$ while experimentally $T_{N}=20K$ we can deduce $J\sim 20$$13\sim 1.5K$. This result of a dominant first-neighbor antiferromagnetic interaction is further supported by the observation of a small (1.8°) deviation from perfect AF alignment between magnetic moments in adjacent layers [37], which may arise from weak Dzyaloshinskii-Moriya interactions (DMI).
@@
GdCu2Ge2 crystallizes in a tetragonal structure [38], with the Gd sublattice adopting a body-centered tetragonal arrangement. Due to the strong elongation of the unit cell ($c/a\gg 1$), considering only the four first-nearest neighbors (NN) would result in a quasi-two-dimensional system, leading to a broadened magnetic transition near $T_{N}$. The four NNN lie in the same crystallographic plane as the NN, preserving the 2D character. To recover the three-dimensional connectivity required to reproduce the experimental specific heat, at least the eight third-nearest neighbors must be included. This implies that a minimum viable model must account for both the first and third neighbor interactions, yielding a total coordination number $z_{0}=12$. As discussed earlier, such a model resembles other lattices with $z=12$, such as FCC(12). In panel (b), we show that the specific heat is reasonably described by the GdCu2Ge2(12) lattice. However, the best agreement with experiment is achieved using the BCC(14) lattice (see panel b and Appendix E). We therefore associate an effective coordination number $z=14$ with this compound. If, to simplify, we consider $J_{1}$ ($z_{1}=4$) equal to $J_{2}$ ($z_{2}=4$) we get the equation
@@
This assumption leads to the possibilities:
@@
  • $|J_{3}/J_{1}|=2.21$ ($z=14$) which we can disregard has it seems unlikely to have a larger strength associated to a more distant bond.
  • $|J_{3}/J_{1}|=0.45$ ($z=14$) appears to be the only reasonable solution.
@@
Without the actual simulation for this particular combination of couplings, we can only estimate the coupling strengths. To do so, we use the mean-field result for the critical temperature (see section V.1).
$\frac{k_{B}T_{C}}{\frac{S(S+1)}{3}}=(z_{1}+z_{2})|J_{1}|+z_{3}|J_{3}|=(8+8\times 0.45)|J_{1}|$
@@
from which we obtain $|J_{1}|=|J_{2}|\sim 0.2$ K and $|J_{3}|\sim 0.09$ K. From these results we can not determine the signs of the coupling, but from the antiferromagnetic nature of the compound, one infers that at least one of these couplings should be antiferromagnetic.
@@
GdPdBi adopts a half-Heusler structure with the Gd sublattice forming a face-centered cubic (FCC) lattice. Fitting with the layered honeycomb lattice (LHc(10), $z=10$) requires an unexpectedly large $J_{2}/J_{1}$ ratio of approximately $6.1$. To validate our approach, we compared our results with inelastic neutron scattering measurements on the related compound GdPtBi ($T_{N}=8.5$ K)[39, 40]. The analysis of GdPtBi[40] from spin-wave theory reported $J_{1}=0.14$ K, $J_{2}=0.27$ K, $J_{3}=0.049$ K, and $K=0.0098$ K, with $|J_{2}|/|J_{1}|>1$ (specifically $|J_{2}|/|J_{1}|\approx 1.93$), demonstrating that second-neighbor interactions dominate in this compound and that the anisotropy is small ($|K|/|J_{1}|\lesssim 0.1$). In GdPtBi, the magnetic structure consists of ferromagnetic (111) planes antiferromagnetically stacked along [111][39, 41], exhibiting low frustration. This structural and magnetic context validated the plausibility of the $|J_{2}|/|J_{1}|\gtrsim 1$ ratio for GdPdBi. But as shown in section VI the specific heat for $|J_{1}|<|J_{2}|$ is almost independent of $J_{2}$ and does not allow us to determine a ratio of interactions.
@@
The crystal structure of GdNiSi3 is an orthorhombic system. Gd atoms forms a bilayered arrangement. Only interaction at the 7th nearest neighbor connects bilayers to provide a tridimensional connectivity ($z_{0}=17$). A good description of the experimental data is obtained considering up to 10th nearest-neighbor ($z=25$) while the best fit is seen considering up to 13th NN ($z=33$). This suggests that long range interactions, well beyond first neighbors, are relevant in this compound. Given the complexity of these interactions we only consider the uniform $J_{\delta}$ case. The GdNiSi3(33) lattice has $T_{C}/{|J|}\sim 145$ while experimentally $T_{N}=22$K so we can deduce $|J|\sim 22$K/145=0.15/145=0.15K. The large amount of couplings needed (up to 13 nearest-neighbors) certainly come from a distribution of interactions. This distribution could be determined with first-principle calculations that considers at least this coordination.
@@
This completes the Gd compounds that we have considered. In a previous work [35], the specific heat of GdCoIn5 has been fitted with essentially the same procedure. The compound has a tetragonal structure. The authors have extrapolated to the thermodynamic limit the specific heat that corresponds to a simple cubic lattice with $z=6$ nearest neighbors with antiferromagnetic interactions, providing an excellent fit of the data.
@@
The Eu sublattice of Eu2Pd3Sn can be thought of as a stacking of corrugated hexagonal layers in the ab plane. The smaller coordination number to have a tridimensional connectivity is $z_{0}=4$. This involves up to the third coordination shell of nearest neighbors (first and second nearest neighbors lie in the same layer). To properly compare the specific heat of Eu2Pd3Sn [43] with a Hamiltonian model we have to consider up to 6th nearest-neighbor in this lattice leading to a coordination $z=14$. The point group at the Eu sites is $C_{1h}$, which allows for the inclusion of the anisotropy term $H_{A}$, as discussed in section II.2. So an anisotropy term with $K/|J_{1}|=-0.4$ was included. With this anisotropy the Eu2Pd3Sn(14) lattice has $T_{C}/{|J_{0}|}\sim 66$. Experimentally $T_{N}=47$K so we can deduce, assuming uniform couplings, $|J|\sim 47$K/66=0.71/66=0.71K.
Finally, the Eu sublattice of EuPdSn2 can be thought of as a stacking of rectangular layers in the ac plane. The smaller coordination number to have a tridimensional connectivity is $z_{0}=8$. This involves up to the third coordination shell of nearest-neighbor interactions. To properly compare the specific heat of EuPdSn2 [31] with a Hamiltonian model we have to consider up to 7th nearest-neighbor in this lattice leading to a coordination $z=22$. The point group at the Eu sites is $C_{2v}$, which also allows for the inclusion of the term $H_{A}$. So an anisotropy term with $K/|J_{0}|=-0.9$ was included. With this anisotropy the EuPdSn2 lattice has $T_{C}/{|J_{0}|}\sim 112$. Experimentally $T_{N}=12$K from which we deduce $|J|\sim 12$K/112=0.11/112=0.11K. The discrepancies near $T_{N}$ likely originate from factors not explicitly included in our model, such as more complex magnetic ordering (e.g., a collinear amplitude-modulated structure) resulting from modulated interactions and crystal-field effects, as discussed in Ref. 14.
@@
While several theoretical curves may appear visually similar to the experimental data, only a single combination of $z$ and $K$ simultaneously minimizes the $L^{2}$ distance and faithfully reproduces the low-temperature specific heat, enabling unambiguous extraction of microscopic parameters.
@@
Figure 8: Experimental specific heat $C/T$ versus $T/T_{N}$ for six 4f7 compounds (dots) and our theoretical fits (lines). $K=0$ for all Gd-based compounds (GdNi3Ga9, GdCu2Ge2, GdPbBi, GdNiSi3). Inset for GdCu2Ge2 show $z$ as a function of $J_{3}/J_{1}$ (black line) and the fitting $z=14$ value (horizontal green line). Experimental data sources: Refs. 36, 37, 44, 38, 29, 31, 43.
@@ VIII Summary
The effective number of neighbors $z$ quantifies the strength of thermal and quantum fluctuations, with larger $z$ values corresponding to reduced fluctuations and specific heat curves closer to mean-field predictions. Crucially, we find that the specific heat is remarkably insensitive to the detailed lattice structure for a given $z$ as long as the couplings do not strongly increase with distance. This enables a universal description of diverse magnetic systems.
@@ Acknowledgements.
Appendix A Mean value and fluctuations of a random variable
@@
Appendix B Distance between specific heat curves
To quantify the similarity between specific heat curves across different lattices, we define a comparison interval that excludes regions dominated by spin-wave excitations and the critical region. Specifically, we discard temperatures below $T/T_{C}=0.25$, where low-energy magnons dominate, and within the critical region $0.95<T/T_{C}<1.05$, where finite-size and scaling effects obscure universal behavior. Above the transition, we restrict our analysis to $T\leq 1.25\,T_{C}$, as above this temperature the specific heat decays rapidly and contributes negligibly to the integrated signal.
The selected interval $[0.25,1.25]$ is divided into 50 equally spaced bins. For each curve, the mean value of the specific heat points falling within each bin is computed. If a curve has no data points in a given bin, that bin is excluded from the comparison for that pair of curves.
The distance between two specific heat curves, $C_{1}(T)$ and $C_{2}(T)$, is then defined as the normalized $L^{2}$ distance over the set of valid bins:
$\ell_{2}=\sqrt{\frac{1}{n_{\text{valid}}}\sum_{i=1}^{n_{\text{valid}}}\left(\langle C_{1}\rangle_{i}-\langle C_{2}\rangle_{i}\right)^{2}},$
where $\langle C_{k}\rangle_{i}$ denotes the mean value of curve $k$ in bin $i$, and $n_{\text{valid}}$ is the number of bins for which both curves have at least one data point. This metric provides a robust measure of dissimilarity between specific heat profiles across lattices, insensitive to small sampling gaps.
Appendix C Similarity of specific heat curves for uniform exchange couplings
Figure 9 presents the $L^{2}$ distance matrix between specific heat curves $C(T)/T$ for 19 non-frustrated lattices with uniform exchange couplings (i.e., $J_{\delta}=J_{1}$), as listed in Table 2. The matrix combines a numerical table with a color heatmap, where off-diagonal entries represent the distance $\ell_{2}$ between pairs of curves, computed over the temperature range $0.25\leq T/T_{C}\leq 1.25$ (see Appendix B).
The color scale ranges from blue (lowest distance, $\ell_{2}\approx 0$) to red (highest distance), with the diagonal entries indicating the clustering structure of lattices by their effective coordination number $z$. The rows and columns are ordered such that lattices with similar $z$ values are grouped together, resulting in a clear block-diagonal structure. This ordering closely follows the sequence of increasing $z$, as defined in Eq. (4).
Lattices with $\ell_{2}<0.03$ are considered to belong to the same cluster, as this threshold corresponds approximately to the visual discriminability limit of curves in Figure 2 and to the estimated numerical uncertainty in our QMC data.
For example, a distinct yellow cluster emerges at $z=26$, comprising the BCC, EuPdSn2, and simple cubic (SC) lattices. Similarly, lattices with $z=18$ (HCP, SC, FCC) and $z=16$ (diamond, EuPdSn2) form well-defined clusters. We also observe clustering between $z=8$ (BCC) and $z=12$ (FCC, HCP), as well as between $z=4$ (diamond, LHc), further supporting the universality of $z$ as the primary determinant of specific heat behavior under uniform couplings.
Figure 9: $L^{2}$ distance matrix between specific heat curves $C(T)/T$ for 19 non-frustrated lattices with uniform exchange couplings ($J_{\delta}=J_{1}$). Off-diagonal elements show the dissimilarity $100\times\ell_{2}$ (i.e., multiplied by 100 for visual clarity), computed over the range $0.25\leq T/T_{C}\leq 1.25$ (see Appendix B). Colors range from blue (low distance, high similarity) to red (high distance, low similarity). Curves with $\ell_{2}<0.03$ are grouped into the same cluster.
Appendix D Similarity of specific heat curves for non-uniform exchange couplings
For systems with non-uniform exchange couplings ($J_{2}\neq J_{1}$), the $L^{2}$ distance $\ell_{2}$ between specific heat curves quantifies deviations from the universality of $z$. Tables 5 and 6 summarize pairwise distances for lattices with $J_{2}<J_{1}$ and $J_{2}>J_{1}$, respectively, computed over the range $0.25\leq T/T_{C}\leq 1.25$.
When $|J_{2}|<|J_{1}|$ (Table 5), curves with the same $z$ remain highly similar ($\ell_{2}<0.06$), consistent with the dominance of $z$ as the controlling parameter. The most notable exception is the layered honeycomb lattice LHc(10), which exhibits slightly larger distances ($\ell_{2}\approx 0.07$$0.11$) relative to cubic lattices, reflecting its distinct low-dimensional topology.
When $|J_{2}|\gg|J_{1}|$ (Table 6), deviations increase significantly for small $z$ ($z=8$, $\ell_{2}\sim 0.1$$0.23$), but diminish as $z$ increases. For $z\geq 14$, distances fall below 0.07, indicating that the universality of $z$ is restored when the effective number of neighbors becomes sufficiently large, even under strong hierarchical couplings.
Table 5: $L^{2}$ distances ($\ell_{2}$) between specific heat curves for lattices with $|J_{2}|<|J_{1}|$.
$z=6$ $\ell_{2}=0.05$
SC(6)  vs  Diamond($J_{2}=-0.080$) 0.05
$z=8$ $\ell_{2}\leq 0.06$
BCC(8)  vs  SC($J_{2}=-0.082$) 0.03
Diamond($J_{2}=-0.155$)  vs  SC($J_{2}=-0.082$ 0.04
BCC(8)  vs  Diamond($J_{2}=-0.155$) 0.06
$z=10$ $\ell_{2}\leq 0.11$
Diamond($J_{2}=-0.236$)  vs  SC($J_{2}=-0.16$) 0.03
Diamond($J_{2}=-0.236$)  vs  BCC($J_{2}=-0.17$) 0.04
SC($J_{2}=-0.16$)  vs  BCC($J_{2}=-0.17$) 0.05
LHc(10)  vs  SC($J_{2}=-0.16$) 0.07
LHc(10)  vs  Diamond($J_{2}=-0.236$) 0.08
LHc(10)  vs  BCC($J_{2}=-0.17$) 0.11
$z=12$ $\ell_{2}\leq 0.04$
SC($J_{2}=-0.25$)  vs  BCC($J_{2}=-0.40$) 0.02
Diamond($J_{2}=-0.333$)  vs  BCC($J_{2}=-0.40$) 0.02
FCC(12)  vs  Diamond($J_{2}=-0.333$) 0.02
Diamond($J_{2}=-0.333$, $z=12$)  vs  SC($J_{2}=-0.25$) 0.03
FCC(12)  vs  BCC($J_{2}=-0.40$) 0.03
FCC(12)  vs  SC($J_{2}=-0.25$) 0.04
$z=14$ $\ell_{2}\leq 0.06$
BCC(14)  vs  FCC($J_{2}=-0.18$) 0.02
SC($J_{2}=-0.35$)  vs  FCC($J_{2}=-0.18$) 0.02
BCC(14)  vs  SC($J_{2}=-0.35$) 0.03
BCC(14)  vs  Diamond($J_{2}=-0.472$) 0.04
Diamond($J_{2}=-0.472$)  vs  FCC($J_{2}=-0.18$) 0.05
Diamond($J_{2}=-0.472$)  vs  SC($J_{2}=-0.35$) 0.06
$z=16$ $\ell_{2}\leq 0.04$
SC($J_{2}=-0.50$)  vs  FCC($J_{2}=-0.40$) 0.02
Diamond(16)  vs  FCC($J_{2}=-0.40$) 0.04
Diamond(16)  vs  SC($J_{2}=-0.50$) 0.04
$z=18$ $\ell_{2}\leq 0.04$
FCC(18)  vs  BCC($J_{1}=J_{2},J_{3}=-0.175$) 0.02
SC(18)  vs  BCC($J_{1}=J_{2},J_{3}=-0.175$) 0.04
Table 6: $L^{2}$ distances ($\ell_{2}$) between specific heat curves for lattices with $|J_{2}|\gg|J_{1}|$.
$z=8$ $\ell_{2}=0.10$$0.23$
BCC($J_{2}=-8$)  vs  FCC($J_{2}=-12.32$) 0.10
Diamond($J_{2}=-0.155$)  vs  BCC($J_{2}=-8$) 0.13
Diamond($J_{2}=-0.155$)  vs  FCC($J_{2}=-12.32$) 0.23
$z=10$ $\ell_{2}=0.09$$0.18$
Diamond($J_{2}=-0.236$)  vs  BCC($J_{2}=-3.82$) 0.09
BCC($J_{2}=-3.82$)  vs  FCC($J_{2}=-6.1$) 0.10
Diamond($J_{2}=-0.236$)  vs  FCC($J_{2}=-6.1$) 0.18
$z=12$ $\ell_{2}=0.08$$0.17$
Diamond($J_{2}=-0.333$)  vs  BCC($J_{2}=-2.27$) 0.08
BCC($J_{2}=-2.27$)  vs  FCC($J_{2}=-4$) 0.11
Diamond($J_{2}=-0.333$)  vs  FCC($J_{2}=-4$) 0.17
$z=14$ $\ell_{2}<0.13$
Diamond($J_{2}=-3.527$)  vs  SC($J_{2}=-5.64$) 0.03
Diamond($J_{2}=-3.527$)  vs  Diamond($J_{2}=-0.472$) 0.04
Diamond($J_{2}=-0.472$)  vs  SC($J_{2}=-5.64$) 0.06
FCC($J_{2}=-2.82$)  vs  SC($J_{2}=-5.64$) 0.07
Diamond($J_{2}=-3.527$)  vs  FCC($J_{2}=-2.82$) 0.08
FCC($J_{2}=-2.82$)  vs  Diamond($J_{2}=-0.472$) 0.13
$z=16$ $\ell_{2}<0.08$
Diamond(16)  vs  SC($J_{2}=-2.5$) 0.04
SC($J_{2}=-2.5$)  vs  FCC($J_{2}=-2$) 0.04
Diamond(16)  vs  FCC($J_{2}=-2$) 0.07
Appendix E Distances between experimental and theoretical specific heat curves
To quantify the quality of the fits presented in Section VII, we compute the $L^{2}$ distance $\ell_{2}$ between each experimental specific heat curve and its theoretical counterparts, using the same temperature range $0.25\leq T/T_{N}\leq 1.25$ (see Appendix B). Table 7 summarizes these distances for all candidate lattices with $\ell_{2}<0.3$, along with relevant anisotropy values where applicable. While the smallest $\ell_{2}$ values identify the closest matches, our selection of the optimal model for each compound prioritizes physical plausibility, especially the reproduction of the low-temperature behavior ($T<0.6\,T_{N}$), which is least affected by uncertainties in phonon subtraction and experimental background correction.
Table 7: $L^{2}$ distances ($\ell_{2}$) between experimental specific heat curves and theoretical predictions for 4f7 compounds. Only models with $\ell_{2}<0.3$ and physically relevant coordination numbers or anisotropy values are shown. For Gd-based compounds, over 20 lattice models were evaluated; for Eu-based compounds, more than 30 combinations of $z$ and $K$ were considered. The selected model for each compound is indicated in bold.
Compound Theoretical lattice $\ell_{2}$
GdNi3Ga9 LHc(4,AF) 0.07
LHc(4) 0.12
LHc(10) 0.15
GdCu2Ge2 FCC(18) 0.07
BCC(14) 0.09
GdCu2Ge2(16) 0.11
GdCu2Ge2(12) 0.17
GdPdBi SC(6) 0.07
BCC(8) 0.07
LHc(10) 0.09
FCC(12) 0.09
FCC(18) 0.19
GdNiSi3 GdNiSi3(33) 0.06
GdNiSi3(25) 0.17
BCC(26) 0.06
SC(18) 0.10
BCC(14) 0.15
Eu2Pd3Sn Eu2Pd3Sn(14, $K=-0.2|J_{1}|$) 0.09
Eu2Pd3Sn(14, $K=-0.4|J_{1}|$) 0.09
Eu2Pd3Sn(14, $K=-0.6|J_{1}|$) 0.10
Eu2Pd3Sn(14, $K=-0.8|J_{1}|$) 0.18
Eu2Pd3Sn(14, $K=0$) 0.20
BCC(26) 0.12
FCC(18) 0.15
Diamond(16) 0.16
Eu2Pd3Sn(20, $K=0$) 0.16
EuPdSn2 EuPdSn2(22, $K=-1.5|J_{1}|$) 0.08
EuPdSn2(22, $K=-0.9|J_{1}|$) 0.10
EuPdSn2(22, $K=-1.2|J_{1}|$) 0.10
EuPdSn2(22, $K=-0.6|J_{1}|$) 0.14
EuPdSn2(22, $K=0$) 0.29
EuPdSn2(16, $K=-0.9|J_{1}|$) 0.10
EuPdSn2(16, $K=-0.6|J_{1}|$) 0.12
EuPdSn2(16, $K=0$) 0.30
BCC(26) 0.22
Diamond(16) 0.30
  • GdNi3Ga9: Best described by the antiferromagnetic layered honeycomb lattice LHc(4,AF), consistent with its magnetic structure described in Section IV and dominant nearest-neighbor interaction $J_{1}$. The small $\ell_{2}$ confirms that long-range interactions are negligible.
  • GdCu2Ge2: Although FCC(18) yields the smallest $\ell_{2}=0.07$, it fails to reproduce the low-temperature specific heat shape. We select BCC(14) ($\ell_{2}=0.09$), which accurately captures the experimental curve in the regime least affected by phonon subtraction artifacts.
  • GdPbBi: The layered honeycomb lattice LHc(10) provides the most consistent description across $T/T_{N}<0.6$, despite SC(6) and BCC(8) having similar $\ell_{2}$. The systematic deviation of SC(6) and BCC(8) below $T_{N}$ renders them less physically plausible for this compound.
  • GdNiSi3: Both GdNiSi3(33) ($\ell_{2}=0.06$) and BCC(26) ($\ell_{2}=0.06$) reproduce the data excellently. The sharp feature near $T/T_{N}\approx 0.8$ is likely an experimental artifact, as no theoretical model reproduces it. We select GdNiSi3(33) because it matches the compound crystal structure.
  • Eu2Pd3Sn: Models with $K=0$ suggest an effective coordination number $z\approx 16$$26$, but only those incorporating axial anisotropy reproduce the suppression of low-temperature entropy. Among the $z=14$ variants, values $K=-0.2|J_{1}|$ to $K=-0.6|J_{1}|$ yield $\ell_{2}<0.10$. We adopt $K=-0.4|J_{1}|$ as the representative value.
  • EuPdSn2: Models with $K=0$ are inconsistent with the experimental data across all coordination numbers. The best fits are obtained for the $z=22$ lattice with axial anisotropy $K=-0.9|J_{1}|$ ($\ell_{2}=0.10$), which accurately reproduces both the low-temperature entropy and the shape of the specific heat peak. Although the $z=16$, $K=-0.9|J_{1}|$ and $z=22$, $K=-1.5|J_{1}|$ models yield equal or slightly lower $\ell_{2}$ values of $0.10$ and $0.08$, respectively, they both underestimate the low-temperature specific heat. The complete failure of all $K=0$ models confirms that axial anisotropy is essential for describing the magnetic thermodynamics of Eu2+ compounds.
These selections underscore that while minimal $\ell_{2}$ provides a useful quantitative guide, the physical consistency to the data remains paramount in identifying the correct magnetic model.